# Anchors Aweigh! Sail for Optimal Unified Multi-Modal Representations

### v2

@@ 1 Introduction
Instead of training separate models for each modality, ImageBind (Girdhar et al., 2023) pairs images with other modalities and projects them into a common image embedding space. Similarly, Zhu et al. (2024) shows that pairing texts with other modalities (LanguageBind) improves cross-modal retrieval performance when language serves as the anchor modality. This approach has inspired various “-Bind” methods tailored to align different modalities for specific domains, such as molecular modeling (Xiao et al., 2024), medical imaging (Gao et al., 2024), brain signals (Yang et al., 2024b), and music selection for videos (Teng et al., 2024). These models commonly use image or text as the anchor embedding due to the abundance of data, with other modalities projected into this anchor representation.
We define the aforementioned approaches as Fixed Anchor Bind (FABind) methods, where the embedding space of the primary anchor modality remains fixed during the alignment process. Many “-Bind” approaches maximize mutual information $I({\bf Z}_{1};{\bf Z}_{i})$ between the representation ${\bf Z}_{1}$ of the anchor modality and the representations ${\bf Z}_{i},i\in\{2,\cdots,M\}$ of other modalities. Although these approaches are practically useful and widely adopted in learning unified multi-modal representation, they have limitations, as demonstrated through theory and experiment in this paper.
@@
First, selecting which modality to serve as an anchor is crucial but challenging, as it depends on both embedding quality and task suitability. Common choices like images or text can be suboptimal, especially when no single modality has the dominant information about the source. Second, fixing an anchor can result in loss of semantic information that is well represented only in other modalities. For instance, while text may describe ‘a dog barks loudly,’ sound could reveal mood, and an image could add facial expression, a holistic combination that a fixed alignment might miss. Third, optimizing only for anchor-to-other-modality overlooks information complementarity between non-anchored modalities. These challenges serve as the primary motivation for introducing adaptive anchor alignment as an alternative to FABind. We formally analyze deficiencies of FABind in Section 2.
Figure 1: Fixed anchor binding methods (FABind) rely on a fixed anchor modality, while adaptive anchors (CentroBind) dynamically generate anchoring points from multi-modal batch samples. Colors and shapes represent modalities and semantic information, respectively, with $\mathcal{Z}$ as the unified representation space.
Figure 2: Graphical illustration of adaptive anchor alignment: Adaptive anchors are dynamically generated from current embeddings using an aggregation method (CentroBind) for every batch.
@@
Encoders are then trained to minimize the ensemble of InfoNCE loss (Oord et al., 2018) between the representation of this adaptive anchor and the fixed modality representations. The main intuition is that a desirable anchor should be representative of all modalities, capturing the most comprehensive information, with well-trained encoders producing representations that naturally cluster around this shared centroid, reflecting their underlying semantic alignment.
Our theoretical analysis demonstrates that CentroBind effectively addresses three critical components of multi-modal learning: 1) capturing intra-modal mutual information, 2) learning inter-modal mutual information, and 3) performing multi-modal alignment by maximizing embedding similarity measures. By incorporating these elements, CentroBind outperforms other multi-modal alignment methods, as shown empirically on both synthetic and real-world datasets in retrieval and classification tasks. The proposed approach yields a unified representation space and, in the perspective of (Huh et al., 2024), who contends that multi-modal representations better align as they move toward a platonic representation that captures the semantic information of all modalities simultaneously.
@@ 2.1 Representation learning framework
Boldface upper case letters (e.g., $\mathbf{X}$) denote random vectors, and a realization is denoted by the boldface lower case letters (e.g., ${\boldsymbol{x}}$); For $n\in\mathbb{N}$, $[n]:=\{1,2,\cdots,n\}$; $P_{{\bf X}}$ and $P_{{\bf X},{\bf Y}}$ denote the marginal and the joint distributions of ${\bf X}$ and $({\bf X},{\bf Y})$, respectively.
@@
We note that if $i=l$, the sufficient encoder provides embeddings with maximum intra information, and if $i\neq l$, it gives embeddings with maximum shared information between $i$-th and $l$-th modalities. [^2] With a proper choice of $\mathcal{Z}_{i}$ ensuring $\max_{f:\mathcal{X}_{i}\to\mathcal{Z}_{i}}I(f({\bf X}_{i});{\bf X}_{l})=I({\bf X}_{i};{\bf X}_{l})$, Definition 2.1 says that $f_{i}({\boldsymbol{x}}_{i,j})$ is a sufficient statistic (Polyanskiy & Wu, 2024) of ${\boldsymbol{x}}_{i,j}$ for ${\boldsymbol{x}}_{l,j}$ as the encoding entails no information loss.
@@
$I_{\rm NCE}({\bf U};{\bf V}|\tau)$ $=\mathbb{E}_{P}\left[-\log\frac{e^{\frac{{\bf U}^{\top}{\bf V}}{\tau}}}{e^{\frac{{\bf U}^{\top}{\bf V}}{\tau}}+\sum_{i=1}^{N}e^{\frac{{\bf U}^{\top}{\bf V}_{i}}{\tau}}}\right],$ (2)
@@ 2.3 Analysis of FABind
Proposition 2.2 shows that the FABind encoders $\mathcal{F}^{\rm FB}$ learned with a sufficient anchor embedding can achieve the maximum shared information between the anchor and the other modalities. However, it does not guarantee shared information between non-anchored modalities $I(f_{i}({\bf X}_{i});f_{l}({\bf X}_{l})),~{}i,l\neq 1$, which can also be seen from (4). Proposition 2.3 establishes that an insufficient anchor may lead to a reduction of shared information between the anchor and the other modalities, implying that the performance of FABind may overly depend on the quality of the arbitrarily selected anchor. Next, we enumerate three limitations in FABind:
P1. Over-reliance on a single anchor modality.
@@
P2. Failure to capture intra information.
@@
P3. Absence of shared information among non-anchored modalities.
@@
We identify three key limitations (P1, P2, and P3) of FABind, supported by empirical evidence presented in Section 4.1. In addition to these constraints, the representations generated by FABind may not approximate an ideal standard, such as the “Platonic representation” described by Huh et al. (2024). Achieving an optimal multi-modal representation necessitates the comprehensive integration of all modalities to fully capture their information. However, FABind falls short of this objective. To address these shortcomings, we next introduce an alternative multi-modal representation that fully leverages sample-level information across all modalities.
@@ 3 Alignment using adaptive anchor
Algorithm 1 CentroBind
  Initialize encoders $f_{1}^{(0)},f_{2}^{(0)},\cdots,f_{M}^{(0)}$.
  for $t=0,1,\ldots,t_{\max}$ do
     Sample $B$ from multi-modal datasets $\{\mathcal{D}_{i}\}_{i}$.
     for $i=1,\ldots,M$ do
        Minimize $\mathcal{L}_{\rm CB}(f_{i}^{(t+1)}|\tau)$ in (8)
     end for
  end for
@@ 3.1 CentroBind
Consider $M$ modalities with corresponding encoders $\{f_{i}\}_{i=1}^{M}$. The CentroBind algorithm is presented in Algorithm 1, and a graphical illustration is given in Figure 2. In the following, we describe each step of the algorithm.
@@
$I_{\rm NCE}({\bf A};f_{i}({\bf X}_{i})|\tau)$ $=\frac{-1}{|\mathcal{I}_{B}|}\sum_{k=1}^{|\mathcal{I}_{B}|}\log\frac{\exp(\frac{{\boldsymbol{a}}_{k}^{\top}f_{i}({\boldsymbol{x}}_{i,k})}{\tau})}{\sum_{j\in\mathcal{I}_{B}}\exp(\frac{{\boldsymbol{a}}_{k}^{\top}f_{i}({\boldsymbol{x}}_{i,j})}{\tau})}$ (9a)
$I_{\rm NCE}(f_{i}({\bf X}_{i});{\bf A}|\tau)$ $=\frac{-1}{|\mathcal{I}_{B}|}\sum_{k=1}^{|\mathcal{I}_{B}|}\log\frac{\exp(\frac{{\boldsymbol{a}}_{k}^{\top}f_{i}({\boldsymbol{x}}_{i,k})}{\tau})}{\sum_{j\in\mathcal{I}_{B}}\exp(\frac{f_{i}^{\top}({\boldsymbol{x}}_{i,k}){\boldsymbol{a}}_{j}}{\tau})}.$ (9b)
@@ 3.2 Theoretical analysis of CentroBind
$|\mathcal{I}_{B}|I_{\rm NCE}\left({\bf A};f_{i}({\bf X}_{i})\;\middle|\;{\tau}\right)$
$\geq\sum_{l=1}^{M}I_{\rm NCE}\left(f_{l}({\bf X}_{l}^{\prime});f_{i}({\bf X}_{i})\;\middle|\;\frac{{\tau}M}{|\mathcal{I}_{B}|}\right)-\sum_{k=1}^{|\mathcal{I}_{B}|}\log C_{\mathcal{F},k,i},$ (10)
where
$C_{\mathcal{F},k,i}=\frac{(c_{\mathcal{F},k,i}^{\min}+c_{\mathcal{F},k,i}^{\max})^{2}}{4c_{\mathcal{F},k,i}^{\min}c_{\mathcal{F},k,i}^{\max}}$ (11)
with $g(l,j|k,i):=\exp\left(\frac{|\mathcal{I}_{B}|f_{l}^{\top}({\boldsymbol{x}}_{l,k}^{\prime})f_{i}({\boldsymbol{x}}_{i,j})}{\tau M}\right)$,
$c_{\mathcal{F},k,i}^{\min}=\min_{l\in[M],j\in\mathcal{I}_{B}}g(l,j|k,i),~{}\text{ and }~{}$
$c_{\mathcal{F},k,i}^{\max}=\max_{l\in[M],j\in\mathcal{I}_{B}}g(l,j|k,i).$ (12)
@@
Theorem 3.1 provides a lower bound of $I_{\rm NCE}\left({\bf A};f_{i}({\bf X}_{i})\;\middle|\;{\tau}\right)$ in (9a), which is a part of the CentroBind objective function $\mathcal{L}_{\rm CB}(f_{i}|\tau)$. Thus CentroBind minimizes a lower bound (3.1) that consists of two terms, $\sum_{l=1}^{M}I_{\rm NCE}\left(f_{l}({\bf X}_{l}^{\prime});f_{i}({\bf X}_{i})\;\middle|\;\frac{{\tau}M}{|\mathcal{I}_{B}|}\right)$ and $-\sum_{k=1}^{|\mathcal{I}_{B}|}\log C_{\mathcal{F},k,i}$. We next provide intuition on why a minimization of the lower bound is justified.
@@
The effect of minimizing $-\sum_{k=1}^{|\mathcal{I}_{B}|}\log C_{\mathcal{F},k,i}$.
We show the effect of growing $C_{\mathcal{F},k,i}$ in terms of cosine similarity score between embeddings. Since $C_{\mathcal{F},k,i}=\frac{1}{4}\left(\sqrt{\gamma}+\sqrt{\frac{1}{\gamma}}\right)^{2}$ with $\gamma=\frac{c_{\mathcal{F},k,i}^{\max}}{c_{\mathcal{F},k,i}^{\min}}\geq 1$, maximizing $C_{\mathcal{F},k,i}$ is equivalent to simultaneously maximizing $c_{\mathcal{F},k,i}^{\max}$ and minimizing $c_{\mathcal{F},k,i}^{\min}$. For ease of analysis, we assume that the encoders are reasonably well-trained. Then, since a positive pair of embeddings normally yields higher similarity score, $c_{\mathcal{F},k,i}^{\max}$ is attained by choosing $l=i$ and $j=k$ in (3.1) as such choices make ${\boldsymbol{x}}_{l,k}^{\prime}$ be positive pair with ${\boldsymbol{x}}_{i,j}$. Thus, $c_{\mathcal{F},k,i}^{\max}$ is roughly proportional to the similarity score of a positive pair of embeddings. Conversely, $c_{\mathcal{F},k,i}^{\min}$ corresponds to the similarity scores of negative pairs, which tend to be low. Hence, minimizing $-\sum_{k=1}^{|\mathcal{I}_{B}|}\log C_{\mathcal{F},k,i}$ enhances the similarity scores for positive pairs and reduces those for negative pairs, improving the overall multi-modal alignment.
@@ 4.1 Experiments with synthetic datasets
Synthetic datasets.
We utilize a Gaussian mixture model (GMM) (Bishop & Nasrabadi, 2006) to generate synthetic multi-modal datasets. Let ${\bf X}_{i}$ denote the $i$-th modality dataset, with modal qualities arranged in ascending order based on their inherent correlation with the latent variable. Specifically, among $M$ modalities, ${\bf X}_{1}$ represents the least correlated modality, while ${\bf X}_{M}$ represents the most correlated. The class label, ${\bf Y}$, is assigned according to the GMM component identifier. A detailed description of the dataset generation process is provided in Appendix C.1.
(a) Comparison of the bind methods.
(b) Random vs. pretrained backbone.
Figure 3: Accuracy as a measure of the representation space quality. Abbreviation: ${\bf X}_{i}$-B or CB: applying FABind with anchor ${\bf X}_{i}$ or applying CentroBind; acc$({\bf Z}_{i})$ or acc(All): accuracy of ${\bf Z}_{i}$ or of concatenated embeddings $({\bf Z}_{1},\cdots,{\bf Z}_{M})$; (rnd): if random backbones are used.
Experiment results.
Figure 3 shows the classification accuracies with a synthetic dataset of $M=4$ modalities. To obtain the results in Figure 3(a), we initialize pretrained backbones for each modality, apply FABind (${\bf X}_{i}$-B) with anchor ${\bf X}_{i}$ or CentroBind (CB), and evaluate accuracy (acc(${\bf Z}_{i}$)) with embeddings from $i$-th modality. We provide acc(${\bf Z}_{i}$), without any binding, for a reference. Figure 3(a) verifies our analysis of FABind (Section 2.3) and CentroBind (Section 3.2): (1) the comparison between ${\bf X}_{1}$-B and ${\bf X}_{4}$-B shows the importance of choosing an anchor modality; (2) the comparison between acc(${\bf Z}_{4}$) and ${\bf X}_{1}$-B: acc(${\bf Z}_{4}$) shows a performance deterioration by FABind , demonstrating the impact on performance of FABind failure to capture intra information; (3) the proposed CB consistently outperforms FABind, indicating that CB successfully captures elements that FABind overlooks, including intra information and shared information among non-anchored modalities.
Figure 3(b) includes accuracies of FABind and CB with random backbones. Similarly, CB outperforms all baselines. Somewhat surprisingly, CB with random backbones (green curves) also performs better than FABind with pretrained backbones (red curves). This further supports our analysis that CentroBind is robust to backbone quality as it optimizes intra and shared information, whereas FABind is sensitive to the backbone quality. Overall, these empirical results validate our analysis of FABind and CentroBind. For clarity, we summarize the final accuracies in Table 3. We provide experimental results with $M=6,8$ in Appendix C.1. With the larger number of modalities, CB still outperforms the baselines.
Embedding visualization.
In addition, we visualize the embeddings learned by FABind and CentroBind using UMAP (McInnes et al., 2018) in Figure 4 (for details and visualizations using t-SNE (Van der Maaten & Hinton, 2008), see Appendix C.1). Figure 4 shows that CentroBind embeddings are better clustered, whereas FABind embeddings appear more scattered, implying that CentroBind yields better embeddings compared to FABind.
Stability.
We empirically examine the stability and convergence speed of both CentroBind and FABind. In Figure 7, we plot the training loss curves, which show similar behavior, suggesting that the adaptive anchor does not introduce issues related to convergence or stability. Detailed discussions can be found in Appendix C.1.
(a) FABind.
(b) CentroBind.
Figure 4: Representation visualization via UMAP.
4.2 Experiments with real-world dataset
In this section, we provide experiment results with a real-world dataset. We compare CentroBind, FABind anchored at text modality, UniBind (Lyu et al., 2024), AudioCLIP (Guzhov et al., 2022), and ViT-Lens (Lei et al., 2024) (see implementation detail in Appendix C.2). We utilize the MUStARD (Castro et al., 2019) and DreamBooth dataset (Ruiz et al., 2023). [^4] For experiments on the DreamBooth dataset, please refer to Appendix C.3. The MUStARD dataset consists of 690 video, audio and text for sarcasm detection with labels such as sarcasm indicators and speaker names. For the backbones in FABind and CentroBind , we use the pretrained VideoMAE (Tong et al., 2022) for video data, the pretrained WaveLM (Chen et al., 2022) for audio data, and the pretrained BERT (Devlin et al., 2019) for text data. A detailed description of the training setting is provided in Appendix C.2.

### v3

@@ 1 Introduction
Instead of training separate models for each modality, ImageBind (Girdhar et al., 2023) pairs images with other modalities and projects them into a common image embedding space. Similarly, (Zhu et al., 2024) shows that pairing texts with other modalities (LanguageBind) improves cross-modal retrieval performance when language serves as the anchor modality. This approach has inspired various “-Bind” methods tailored to align different modalities for specific domains, such as molecular modeling (Xiao et al., 2024), medical imaging (Gao et al., 2024), brain signals (Yang et al., 2024b), and music selection for videos (Teng et al., 2024). These models commonly use image or text as the anchor embedding due to the abundance of data, with other modalities projected into this anchor representation.
We refer to these approaches as Fixed Anchor Bind (FABind) methods, where the primary anchor modality’s embedding space is held fixed while aligning others. Many “-Bind” methods maximize mutual information $I({\bf Z}_{1};{\bf Z}_{i})$ between the anchor representation ${\bf Z}_{1}$ and each non-anchor ${\bf Z}_{i}$ for $i\in{2,\ldots,M}$. Despite their practical appeal for unified multimodal representations, we show, both theoretically and empirically, that they have important limitations.
Figure 1: Limitations of fixed-anchor binding (FABIND) and remedy via CentroBind (CB). Shapes denote semantic classes (e.g., dog, cat, horse) and colors denote modalities (image, text, audio). P1–Over-reliance on anchor: a poor or poorly chosen anchor yields misalignment across modalities. P2–Loss of intra-modal information: anchoring suppresses modality-specific cues present only in non-anchored views. P3–Loss of shared information among non-anchors: optimizing only anchor $\leftrightarrow$ others ignores correlations between non-anchored modalities. CB addresses all three by computing adaptive, batch-wise anchors (centroids) from the available modalities and aligning each modality to this shared anchor, preserving intra-information and non-anchor shared-information while improving overall alignment.
@@
Figure 2: Graphical illustration of adaptive anchor alignment: Adaptive anchors are dynamically generated from current embeddings using an aggregation method for every batch.
As illustrated in Figure 1, we discover three limitation of FABind. P1–over-reliance on anchor: the best anchor choice depends on embedding quality and task, and common defaults (image or text) can be suboptimal when no single modality dominates. P2–loss of intra-information: a fixed anchor can discard semantics captured primarily by other modalities (e.g., sound revealing mood, images conveying expression beyond text like “a dog barks loudly”). P3–loss of shared-information: optimizing only anchor–other pairs ignores complementarities among non-anchor modalities. These issues motivate adaptive anchor alignment as an alternative to FABind; we formalize FABind’s deficiencies in Section 2.
@@
Our theoretical analysis demonstrates that CentroBind effectively addresses three critical components of multi-modal learning: 1) capturing intra-modal mutual information, 2) learning inter-modal mutual information, and 3) performing multi-modal alignment by maximizing embedding similarity measures. By incorporating these elements, CentroBind outperforms other multi-modal alignment methods, as shown empirically on both synthetic and real-world datasets in various downstream tasks. The proposed approach yields a unified representation space and, in the perspective of (Huh et al., 2024), who contends that multi-modal representations better align as they move toward a platonic representation that captures the semantic information of all modalities simultaneously.
@@ 2.1 Representation learning framework
Boldface upper case letters (e.g., $\mathbf{X}$) denote random vectors, and a realization is denoted by the boldface lower case letters (e.g., $\boldsymbol{x}$}); For $n\in\mathbb{N}$, $n]:=\{1,2,\cdots,n\}$; $P_{{\bf X}}$ and $P_{{\bf X},{\bf Y}}$ denote the marginal and the joint distributions of $\bf X$} and $({\bf X},{\bf Y})$, respectively.
@@
We note that if $i=l$, the sufficient encoder provides embeddings with maximum intra information, and if $i\neq l$, it gives embeddings with maximum shared information between $i$-th and $l$-th modalities. [^2] With $\mathcal{Z}_{i}$ such that $\max_{f:\mathcal{X}_{i}\to\mathcal{Z}_{i}}I(f({\bf X}_{i});{\bf X}_{l})=I({\bf X}_{i};{\bf X}_{l})$, Definition 2.1 says that $f_{i}({\boldsymbol{x}}_{i,j})$ is a sufficient statistic (Polyanskiy & Wu, 2024) of ${\boldsymbol{x}}_{i,j}$ for ${\boldsymbol{x}}_{l,j}$ as the encoding entails no information loss.
@@
$I_{\rm NCE}({\bf U};{\bf V}|\tau)$ $=\mathbb{E}_{P}\left[-\log\frac{\exp({\bf U}^{\top}{\bf V}/\tau)}{\exp({\bf U}^{\top}{\bf V}/\tau)+\sum_{i=1}^{N}\exp({\bf U}^{\top}{\bf V}_{i}/\tau)}\right],$ (2)
@@ 2.3 Analysis of FABind
Proposition 2.2 shows that the FABind encoders $\mathcal{F}^{\rm FB}$ learned with a sufficient anchor embedding can achieve the maximum shared information between the anchor and the other modalities. However, it does not guarantee shared information between non-anchored modalities $I(f_{i}({\bf X}_{i});f_{l}({\bf X}_{l})),~i,l\neq 1$, which can also be seen from (4). Proposition 2.3 establishes that an insufficient anchor may lead to a reduction of shared information between the anchor and the other modalities, implying that the performance of FABind may overly depend on the quality of the arbitrarily selected anchor. Next, we enumerate three limitations in FABind, as illustrated in Figure 1:
P1–Over-reliance on a single anchor modality.
@@
P2–Failure to capture intra information.
@@
P3–Absence of shared information among non-anchored modalities.
@@
We identify three key limitations (P1, P2, and P3) of FABind, supported by empirical evidence presented in Section 4.1. In addition to these constraints, the representations generated by FABind may not approximate an ideal standard, such as the “Platonic representation” described in (Huh et al., 2024). Achieving an optimal multi-modal representation necessitates the comprehensive integration of all modalities to fully capture their information. However, FABind falls short of this objective. To address these shortcomings, we next introduce an alternative multi-modal representation that fully leverages sample-level information across all modalities.
@@ 3.1 CentroBind
Consider $M$ modalities with corresponding encoders $\{f_{i}\}_{i=1}^{M}$. The CentroBind algorithm is presented in Algorithm 1 in Appendix C, and a graphical illustration is given in Figure 2. In the following, we describe each step of CentroBind.
@@
$I_{\rm NCE}({\bf A};f_{i}({\bf X}_{i})|\tau)$ $=\frac{-1}{|\mathcal{I}_{B}|}\sum_{k=1}^{|\mathcal{I}_{B}|}\log\frac{\exp({\boldsymbol{a}}_{k}^{\top}f_{i}({\boldsymbol{x}}_{i,k})/\tau)}{\sum_{j\in\mathcal{I}_{B}}\exp({\boldsymbol{a}}_{k}^{\top}f_{i}({\boldsymbol{x}}_{i,j})/\tau)}$ (9a)
$I_{\rm NCE}(f_{i}({\bf X}_{i});{\bf A}|\tau)$ $=\frac{-1}{|\mathcal{I}_{B}|}\sum_{k=1}^{|\mathcal{I}_{B}|}\log\frac{\exp({\boldsymbol{a}}_{k}^{\top}f_{i}({\boldsymbol{x}}_{i,k})/\tau)}{\sum_{j\in\mathcal{I}_{B}}\exp(f_{i}^{\top}({\boldsymbol{x}}_{i,k}){\boldsymbol{a}}_{j}/\tau)}.$ (9b)
@@ 3.2 Theoretical analysis of CentroBind
$|\mathcal{I}_{B}|I_{\rm NCE}\left({\bf A};f_{i}({\bf X}_{i})\;\middle|\;{\tau}\right)\geq\sum_{l=1}^{M}I_{\rm NCE}\left(f_{l}({\bf X}_{l}^{\prime});f_{i}({\bf X}_{i})\;\middle|\;\frac{{\tau}M}{|\mathcal{I}_{B}|}\right)-\sum_{k=1}^{|\mathcal{I}_{B}|}\log C_{\mathcal{F},k,i},$ (10)
where $C_{\mathcal{F},k,i}=\frac{(c_{\mathcal{F},k,i}^{\min}+c_{\mathcal{F},k,i}^{\max})^{2}}{4c_{\mathcal{F},k,i}^{\min}c_{\mathcal{F},k,i}^{\max}}$ with $g(l,j|k,i):=\exp\left(\frac{|\mathcal{I}_{B}|f_{l}^{\top}({\boldsymbol{x}}_{l,k}^{\prime})f_{i}({\boldsymbol{x}}_{i,j})}{\tau M}\right)$,
[…]
@@
Theorem 3.1 provides a lower bound of $I_{\rm NCE}\left({\bf A};f_{i}({\bf X}_{i})\;\middle|\;{\tau}\right)$ in (9a), which is a part of the CentroBind objective function $\mathcal{L}_{\rm CB}(f_{i}|\tau)$. Thus CentroBind minimizes a lower bound (10) that consists of two terms, $\sum_{l=1}^{M}I_{\rm NCE}\left(f_{l}({\bf X}_{l}^{\prime});f_{i}({\bf X}_{i})\;\middle|\;\frac{{\tau}M}{|\mathcal{I}_{B}|}\right)$ and $\sum_{k=1}^{|\mathcal{I}_{B}|}\log C_{\mathcal{F},k,i}$. We next provide intuition on why a minimization of the lower bound is justified.
@@
The effect of minimizing $\sum_{k=1}^{|\mathcal{I}_{B}|}\log C_{\mathcal{F},k,i}$.
We show the effect of growing $C_{\mathcal{F},k,i}$ in terms of cosine similarity score between embeddings. Since $C_{\mathcal{F},k,i}=\frac{1}{4}\left(\sqrt{\gamma}+\sqrt{\frac{1}{\gamma}}\right)^{2}$ with $\gamma=\frac{c_{\mathcal{F},k,i}^{\max}}{c_{\mathcal{F},k,i}^{\min}}\geq 1$, maximizing $C_{\mathcal{F},k,i}$ is equivalent to simultaneously maximizing $c_{\mathcal{F},k,i}^{\max}$ and minimizing $c_{\mathcal{F},k,i}^{\min}$. For ease of analysis, we assume that the encoders are reasonably well-trained. Then, since a positive pair of embeddings normally yields higher similarity score, $c_{\mathcal{F},k,i}^{\max}$ is attained by choosing $l=i$ and $j=k$ in (11) as such choices make ${\boldsymbol{x}}_{l,k}^{\prime}$ be positive pair with ${\boldsymbol{x}}_{i,j}$. Thus, $c_{\mathcal{F},k,i}^{\max}$ is roughly proportional to the similarity score of a positive pair of embeddings. Conversely, $c_{\mathcal{F},k,i}^{\min}$ corresponds to the similarity scores of negative pairs, which tend to be low. Hence, minimizing $\sum_{k=1}^{|\mathcal{I}_{B}|}\log C_{\mathcal{F},k,i}$ enhances the similarity scores for positive pairs and reduces those for negative pairs, improving the overall multi-modal alignment.
@@ 4 Experiment
To thoroughly evaluate the effectiveness of the proposed method, we design two sets of experiments: (1) experiments on synthetically generated datasets and (2) experiments on real-world datasets spanning diverse modality domains. The synthetic datasets allow controlled testing of extreme scenarios, such as varying numbers of modalities, modality imbalance, and backbone quality. The real-world experiments demonstrate that the proposed method generalizes well across datasets of different scales: DreamBooth (Ruiz et al., 2023) ($\sim$180 images across 30 subjects), MUStARD (Castro et al., 2019) ($\sim$690 labeled clips), AVE (Tian et al., 2018) ($\sim$4,143 videos across 28 event categories), UR-FUNNY (Hasan et al., 2019) ($\sim$1,866 TED videos from 1,741 speakers, totaling $\sim$90 hours), and AudioSet (Gemmeke et al., 2017) ($\sim$2 million human-labeled 10-second clips across 632 event classes). We compare CentroBind, FABind (anchored at the text modality for MUStARD and at the image modality for the other datasets), UniBind (Lyu et al., 2024), AudioCLIP (Guzhov et al., 2022), and ViT-Lens (Lei et al., 2024).
@@ 4.1 Experiments with synthetic datasets
We conduct controlled experiments on synthetic datasets generated from a Gaussian mixture model (Bishop & Nasrabadi, 2006). Appendix C.1 provides complete details on data generation, experimental settings, and results. These experiments evaluate whether CentroBind resolves the limitations in Section 2.3 across scenarios with varying numbers of modalities, modality imbalance, and encoder quality. Across all configurations, the synthetic results show that CentroBind overcomes the limitations of FABind. Additional analyses (embedding visualizations, stability studies, temperature tuning, and complexity) are also included in Appendix C.1.
4.2 Experiments with real-world datasets
We evaluate and compare the performance of CentroBind and baseline methods on real-world datasets. See Appendix C.2 for detailed discussion on baselines and experimental results.